OpenStudy (deadshot):

What is the simplified form of the quantity 4 z squared minus 16z plus 15 over the quantity 2 z squared plus z minus 15 ?

OpenStudy (anonymous):

\[\frac{ 4z^2-16z+15 }{ 2z^2+z-15 }\] is it yours?

OpenStudy (anonymous):

since Calisto here, let him help. hihi... he is much more better than me

OpenStudy (callisto):

@DeadShot Are you here? Is the expression given by Hoa right?

OpenStudy (callisto):

Suppose the expression is correct. To simplify it, we need to factorize both the numerator and denominator. Consider the numerator, \[4z^2-16z+15\]\[=4z^2-10z-6z+15\]Take out the common factor for the first two terms and last two terms: \[=2z(2z-5)-3(2z-5)\]Take out the common factor again, we'll get \[=(2z-3)(2z-5)\] Now, try to factorize the denominator.

OpenStudy (deadshot):

\[So-the- answer -is\frac{ (2z-3) }{ (z+3) } \]right?